# The Nonlinear Filter Model of Stream Cipher Redivivus

### v3

@@ 3 Construction from Maiorana-McFarland Bent Functions
Suppose $n=2m$ is even. From Theorem 1 we have ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m})\geq{\sf AI}({\sf Maj}_{m})$. Since ${\sf AI}({\sf Maj}_{m})=\lceil m/2\rceil$, we have the result. For odd $n=2m+1$, from Proposition 2, we have ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}\geq{\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$, which shows the result.
We computed the algebraic immunity of $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{n}$ for small values of $n$ and observed that for $n\geq 4$, ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{n})=1+\lfloor n/4\rfloor$. This suggests that the lower bound in Proposition 4 is exact if $n\not\equiv 0\bmod 4$ and if $n\equiv 0\bmod 4$, the actual algebraic immunity is one more than the lower bound. As mentioned in Section 2, for any Boolean function $f$, ${\sf FAI}(f)\geq 1+{\sf AI}(f)$. Our experiments suggest that for $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{n}$, the fast algebraic immunity is actually one more than the algebraic immunity.
So $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{n}$ does not have the best possible algebraic immunity. On the other hand, it can be implemented using $O(m)$ gates (as we show in Section 5). So it is possible to increase the number of variables to achieve the desired level of algebraic resistance without increasing the circuit size too much.
4 Concrete Proposals
From Proposition 2 we have ${\sf MM}_{2m+1}$ is balanced, and the cryptographic resistance provided by ${\sf MM}_{2m+1}$ is very similar to that provided by ${\sf MM}_{2m}$. In terms of implementation efficiency, ${\sf MM}_{2m+1}$ requires one extra XOR gate in addition to the circuit to implement ${\sf MM}_{2m}$. An LFSR is a linear map, so the all zero-state will lead to all-zero output of the LFSR. If ${\sf MM}_{2m+1}$ maps the all-zero string to 0, then the keystream generated from the all zero-state of the LFSR will be the all-zero keystream. To avoid this scenario, we define the function $f_{2m+1}:\{0,1\}^{2m+1}\rightarrow\{0,1\}$ as follows.
$f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ $=$ $\left\{\begin{array}[]{ll}{\sf MM}_{2m+1}(W,\mathbf{X},\mathbf{Y})&\mbox{if }{\sf MM}_{2m+1}(\mathbf{0}_{2m+1})=1,\\ 1\oplus{\sf MM}_{2m+1}(W,\mathbf{X},\mathbf{Y})&\mbox{if }{\sf MM}_{2m+1}(\mathbf{0}_{2m+1})=0.\end{array}\right.$ (7)
The cryptographic properties of $f_{2m+1}$ are exactly the same as those of ${\sf MM}_{2m+1}$, with the only exception being that $f_{2m+1}$ maps $\mathbf{0}_{2m+1}$ to $1$, while ${\sf MM}_{2m+1}$ may or may not do so. We construct $f_{2m+1}$ from $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}$, i.e. where $h$ is chosen to be ${\sf Maj}_{m}$ and $\pi$ is chosen to be ${\sf id}$.
We revisit the filter generator model where the state of an LFSR of length $L$ is filtered using a Boolean function. By $\mathcal{S}(L,m)$, we denote the class of stream ciphers obtained from an LFSR of length $L$ with a primitive connection polynomial, and with $f_{2m+1}$ as the filtering function. For the $\kappa$-bit security level, we assume that the stream cipher supports a $\kappa$-bit secret key and a $\kappa$-bit initialisation vector (IV), and so $L\geq 2\kappa$. Suppose that the state $(s_{L-1},\ldots,s_{0})$ of the LFSR leads to the next state $(s_{L},\ldots,s_{1})$, i.e. the state is right shifted once and the next bit $s_{L}$ becomes the leftmost bit. The initial state of the LFSR is constructed as $(k_{\kappa-1},\ldots,k_{0},v_{\kappa-1},\ldots,v_{0},b_{1},\ldots,b_{L-2\kappa})$, where $(k_{\kappa-1},\ldots,k_{0})$ is the key, $(v_{\kappa-1},\ldots,v_{0})$ is the IV, and $(b_{1},\ldots,b_{L-2\kappa})\neq\mathbf{0}_{L-2\kappa}$ is an arbitrary constant bit string.
We next specify the tap positions, i.e. which state bits of the LFSR are fed into the filtering function. From the state $(s_{L-1},\ldots,s_{0})$ of the LFSR, the value of $W$ in $f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ is the state bit $s_{L-\kappa+m}$, the values of the variables in $\mathbf{X}$ are the state bits $s_{L-\kappa+m-1},\ldots,s_{L-\kappa}$, and the values of the variables in $\mathbf{Y}$ are the state bits $s_{L-2\kappa+m-1},\ldots,s_{L-2\kappa}$. So if we write the state vector $\mathbf{s}=(s_{L-1},\ldots,s_{0})$ as $\mathbf{s}_{1}$, $\mathbf{s}_{2}$ and $\mathbf{s}_{3}$, where $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ are $\kappa$-bit strings and $\mathbf{s}_{3}$ is a $(L-2\kappa)$-bit string, then the tap positions for $W$ and $\mathbf{X}$ are the rightmost $m+1$ bits of $\mathbf{s}_{1}$, and the tap positions for $\mathbf{Y}$ are the rightmost tap positions for $\mathbf{s}_{2}$. Note that the tap positions are divided into two clusters, the first cluster consisting of $m+1$ consecutive tap positions, the second cluster consisting of $m$ tap positions, and there is a gap of $\kappa-m$ positions between the two clusters. Our choice of tap positions does not follow the recommendation given in the literature [27, 28, 33, 46, 58]. Nonetheless, we later argue that the design can resist state guessing attacks. This is due to the fact that we use Boolean functions on many more variables than what has been proposed in the literature. Let $M$ be the length of the interval between the first tap position for $\mathbf{X}$ and the last tap position for $\mathbf{Y}$, including the end points. Then $M=\kappa+m$. The quantity $M$ will be later required to quantify resistance to inversion attacks [27].
We assume that the stream cipher goes through an initialisation phase of $2\kappa$ steps, where no keystream bit is produced, and instead the keystream bit is fed back by XORing it with the next bit of the LFSR. So after the initialisation phase the leftmost $2\kappa$ bits of the LFSR are complex nonlinear functions of the key and the IV. Keystream generation starts after the initialisation phase.
@@ 4 Concrete Proposals
The parameters $L$ and $m$ are the design parameters of the stream cipher. The goal is to choose these parameters so as to resist well known classes of known plaintext attacks which result in recovering the key. We consider two basic attack parameters, namely the number $N$ of keystream bits that is required for a successful attack, and the time complexity $T$ of the attack. To ensure security at level $\kappa$ any attack should require $T>2^{\kappa}$. The condition $L\geq 2\kappa$ ensures that the size of the state is at least twice the size of the key. This prevents certain (theoretical) time/memory trade-off attacks [34]. We assume that at most $2^{B}$ keystream bits are generated from a single key and IV pair. So $N$ is at most $2^{B}$, as otherwise the attack cannot be mounted. Further, a basic condition is that $N\leq T$, i.e. some operation is performed on each keystream bit that is required. So for $\kappa$-bit security using $2^{B}$ keystream bits, we have $B\leq\kappa$. In the analysis below, our goal is to determine values of $L$ and $m$ such that the stream cipher is not vulnerable to the known types of attacks. For this concrete analysis, we ignore small constants appearing in the expressions for $N$ and $T$ (i.e. we consider these constants to be 1) and focus only on the exponential components of these quantities.
@@
The degree of $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}$ is $d=2^{\lfloor\log_{2}m\rfloor}$. If $L$ is a prime, then the linear complexity of the generated keystream sequence is (${L\choose d}$ (see [48]) and so at least these many keystream bits are required to determine the linear complexity. So if
@@
These are guess-then-determine attacks. The idea is to guess some bits and then use the obtained keystream to verify the guess. The first attack of this type was described in [27] and is called the inversion attack. The inversion attack is specifically applicable to filter functions of the form $X\oplus g(\mathbf{Z})$. Since the filter function that we use is indeed of this form, we need to consider the resistance of the stream cipher to the inversion attack. Consider the bit sequence generated by the LFSR to be $s_{0},s_{1},s_{2},\ldots$ and the corresponding keystream bits be $k_{0},k_{1},k_{2},\ldots$. Due to the placement of the tap positions, the $i$-th keystream bit $k_{i}$ is determined by the bits $s_{i},s_{i-1},\ldots,s_{i-M}$ (see above for the definition of $M$). Further, we have $k_{i}=s_{i}\oplus g(s_{i-1},\ldots,s_{i-M})$, where $g$ is defined from ${\sf MM}_{2m}$ as a function of $M$ variables which is degenerate on the $\kappa-m$ variables. So we have $s_{i}=k_{i}\oplus g(s_{i-1},\ldots,s_{i-M})$, i.e. the $i$-th state bit is determined by the known keystream bit $k_{i}$ and the values of the bits $s_{i-1},\ldots,s_{i-M}$. Once $s_{i}$ is obtained, $s_{i+1}$ is obtained as $s_{i+1}=k_{i+1}\oplus g(s_{i},\ldots,s_{i-M+1})$, similarly, from $s_{i+1}$, one obtains $s_{i+2}$ and so on. In other words, if for some $i$, the values of $s_{i-1},\ldots,s_{i-M}$ are known, then the LFSR sequence can be computed in the forward direction. The attack proceeds by guessing the values of $s_{i-1},\ldots,s_{i-M}$, computing the LFSR sequence for $L$ steps to determine the state of the LFSR and then using this state to generate a keystream, if the generated keystream matches with the obtained keystream, then the guess is correct. Since there are $M$ bits whose values have to be guessed, the method requires $2^{M}$ guesses and about $Lm$ operations for each guess. For ${\sf S}(L,m)$, the value of $M$ is $\kappa+m$. So the inversion attack requires more than $2^{\kappa+m}$ operations on ${\sf S}(L,m)$. This proves the resistance of ${\sf S}(L,m)$ to the inversion attack. Note that the inversion attack is prevented due to $M$ being sufficiently large. This was one of the countermeasures to inversion attack that was already proposed in [27].
The inversion attack was later extended to the generalised inversion attack, the filter state guessing attack, and the generalised filter state guessing attack [27, 28, 33, 46, 58]. As one of the countermeasures to these attacks, some design criteria, such as full positive difference set, for choosing tap positions have been suggested. Choosing the tap positions to be consecutive has been mentioned to be the worst possible choice [33]. Our choice of tap positions consists of two clusters where the tap positions in each cluster are consecutive. So based on the recommendations in the literature, our choice of tap positions is close to the worst possible design choice. Nonetheless, we argue that $\mathcal{S}(L,m)$ resists all kinds of state guessing attacks. The reason is that we use a filtering function on many more variables than has been considered in the literature. All the recommendations in the literature implicitly assume that the number of variables of the filtering function is much smaller than the target security level $\kappa$. Since we use filtering functions with number of variables close to $\kappa$, the recommendations on tap positions in the literature do not apply to our case.
Let $n$ be the number of variables of the filtering function. At any point of time, each of the input bits to the filtering function can be written as a linear function of the $L$ initial state bits of the LFSR. So knowing the values of the $n$ input bits at any point of time provide $n$ linear equations in $L$ variables. If the values of the input bits at $c$ points of time are known, where $nc\geq L$, then one obtains a system of $L$ linear equations in $L$ variables and hence can solve this system to obtain the initial state of the LFSR. Of course, one does not know the input bits to the filtering function at any point of time. There are $2^{n}$ possible values of these input bits. The keystream bits $k_{1},k_{2},\ldots$, i.e. the outputs of the filtering function are known. Since the filtering function is balanced, observing one bit of the keystream, say $k_{i}$, rules out half of the possible values of the corresponding input bits to the filtering function, i.e. there are $2^{n-1}$ possible values of the input bits that give rise to the observed keystream bit. So guessing one of these $2^{n-1}$ values provides $n$ linear equations in $L$ variables. An input bit to the filtering function at time point $i$ may also be an input bit to the filtering function at a few other time points. For $\mathcal{S}(L,m)$ the two sets of input bits to the filtering function at two successive time points $i$ and $i+1$ have $12m-$ 1 bits in common. Suppose we have guessed the values of the input bits for time point $i$. Then the number of possible values for the input bits at time point $i+1$ is $4$. Observing $k_{i+1}$ and again using the fact that the filtering function is balanced, the number of possible values for the input bits at time point $i+1$ reduces to $2$. So the observed keystream bits $k_{i}$ and $k_{i+1}$ together restrict the number of possibilites of values of the input bits at time point $i$ and at time point $i+1$ to be $2^{n}$. By guessing one of these $2^{n}$ values, we obtain a system of $2n$ linear equations in $L$ variables. Extending this idea, we see that by guessing one of $2^{n+c-1}$ values of the $c$ consecutive values of the input bits at time points $i,i+1,\ldots,i+c-1$, we obtain a system of $nc$ equations in $L$ variables. If $n+c>L$, then we can solve the linear system and obtain the initial state of the LFSR. Solving the linear system requires $L^{3}$ operations. Since a linear system has to be solved for each guess, the complexity of the attack is $2^{n+c-1}L^{3}$. Suppose the following condition holds.
@@
$2^{n+c-1}L^{3}$ $>$ $2^{\kappa}\mbox{ where $c$ is the least positive integer such that $n+c>L$.}$ (9)
@@
Then no form of the state guessing attack succeeds against $\mathcal{S}(L,m)$ at the $\kappa$-bit security level.
@@
In the above analysis, we saw that the tap positions of $\mathcal{S}(L,m)$ (two clusters, with consecutive positions in each cluster) result in the fact that once the values of the input bits to the $i$-th position is guessed, there are only two possible values for the input bits to the $(i+1)$-st position. This is due to the fact of the huge overlap between the sets of input bits in the $i$-th and $(i+1)$-st positions. The recommendations for tap positions in the literature [27, 28, 33, 46, 58] aim to reduce this overlap. A commonly used recommendation is full positive difference set, i.e. the absolute values of the differences between the tap positions should be distinct. For a filtering function with $n$ variables, this recommendation results in the gap between the first and the last tap positions to be more than $21+2+\cdots+(n-1)=n(n-1)/$ 2 (since the successive differences must be distinct). So the size of the LFSR is at least quadratic in $n$. From efficiency considerations, this forces the value of $n$ to be small. By not following this recommendation, we have done away with the condition that $L$ must be at least quadratic in $n$. Instead, by considering the fundamental requirement behind state guessing attacks, we identified (9) as the condition to resist such attacks. This allows use to choose $n$ to be quite large and close to $\kappa$. Of course, this is possible due to the fact that we have an efficient method for implementing the filtering functions on such large values of $n$.
@@
Subsequent works (such as [60, 59, 39, 40]) on fast correlation attacks use vectorial decoding technique along with multiple linear approximations, use of the BKW algorithm [6], and multivariate correlation attack. The attacks are quite complex and there are no simple closed form expressions for the values of $N$ and $T$. The stream ciphers to which these attacks are applied are Grain-128a and Sosemanuk [5], which is another cipher providing 128-bit security. From the above discussion, we already know that the linear bias of $\mathcal{S}(257,59)$ is substantially lower than that of Grain-128a. For Sosemanuk, the best known [39] linear approximation has correlation $2^{-20.84}$, which is far greater than the linear bias $2^{-60}$ for $\mathcal{S}(257,59)$. The very low linear bias of $\mathcal{S}(257,59)$ and in general of the other stream cipher proposals in Table 1 make the attacks in [60, 59, 39, 40] inapplicable at the stated security levels.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $N^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [14]). We assume $\omega=2.8$. From Proposition 4, we conservatively assume that ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=\lceil m/2\rceil$, i.e. the algebraic immunity is equal to its lower bound. Let

### v5

@@ 4.1 Effect of Shifts on Tap Positions
From Theorem 1, we have that any monomial occuring in the ANF of ${\sf Maj}_{m}(\mathbf{U},\mathbf{V})$ has degree more than $\lfloor m/2\rfloor$. So any monomial occurring in the ANF of ${\sf Maj}_{m}(\mathbf{U},\mathbf{V})$ must involve one of the $V_{i}$’s, and similarly, any monomial occurring in the ANF of ${\sf Maj}_{m}(\mathbf{U},\mathbf{W})$ must involve one of the $W_{j}$’s. Since the ANF of ${\sf Maj}_{m}(\mathbf{U},\mathbf{V})$ does not involve any $W_{j}$, and the ANF of ${\sf Maj}_{m}(\mathbf{U},\mathbf{W})$ does not involve any $V_{i}$, there can be no cancellation of terms between the two ANFs.
Proposition 6 assures us that if the overlap between two calls to majority is at most $\lfloor m/2\rfloor$, then there is no cancellation of terms arising from the ANF corresponding to the two calls.
To generate keystream bits, the function $f\circ{\sf proj}$ is applied to $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$. This in turn requires applying the majority function to a subset of the bits of $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$. More precisely, the functions calls ${\sf Maj}(s_{L-1-i_{1}},\ldots,s_{L-1-i_{m}})$ and ${\sf Maj}(s_{L+t-1-i_{1}},\ldots,s_{L+t-1-i_{m}})$ occur corresponding to $\mathbf{s}^{(0)}$ and $\mathbf{s}^{(t)}$ respectively. If $t<\kappa$, then there is a possibility that the inputs to the two calls to majority have an overlap, i.e. some of the inputs are common to both the majority calls. Since majority is a symmetric function, the actual positions where the common inputs occur do not matter. For $t\geq 1$, let
$\nu_{t}$ $=$ #($\#\left(\{L-1-i_{1},\ldots,L-1-i_{m}\}\cap\{L+t-1-i_{1},\ldots,L+t-1-i_{m}\}\right).$ (15)
Note that $\nu_{t}=0$ for $t\geq\kappa$, and for $1\leq t<\kappa$,
$\nu_{t}$ $=$ ${\sf wt}({\sf posX}\wedge({\sf posX}\gg t)),$ (16)
where $\wedge$ denotes the bitwise AND operation, and tmuch-greater-$\gg t$ denotes right shift by $t$ places. Define
$\nu$ $=$ $\max_{1\leq t<\kappa}\nu_{t}=\max_{1\leq t<\kappa}{\sf wt}({\sf posX}\wedge({\sf posX}\gg t)).$ (17)
A design goal is to choose ${\sf posX}$ so as to minimise $\nu$. In particular, if $\nu$ is at most $\lfloor m/2\rfloor$, then from Proposition 6 there is no cancellation of terms arising from the two majority calls. Later we provide concrete values of $\nu$ (see Table 4) for our specific proposals. For all cases, the condition $\nu\leq\lfloor m/2\rfloor$ holds, and so the no cancellation property of terms arising from the ANF of the majority function corresponding to the two calls is ensured.
Overlap of tap positions due to shifts.
Let
$\delta$ $=$ $\max_{1\leq t<2\kappa}{\sf wt}({\sf pos}\wedge({\sf pos}\gg t)).$ (18)
Then $\delta$ is the maximum overlap between the tap positions that can be obtained by shifting the state by any value.
Procedure for choosing the tap positions.
Suppose $\kappa$, $L\geq 2\kappa$, and $m$ are given. We need to determine $i_{1},\ldots,i_{m}$ and $j_{1},\ldots,j_{m}$, or equivalently, to determine ${\sf pos}$. (Recall that the tap position for $W$ is fixed to be $L-2\kappa$.) The leftmost $\kappa$ bits of ${\sf pos}$ is the string ${\sf posX}$. The value of $\nu$ is determined by ${\sf posX}$, while the value of $\delta$ is determined by ${\sf pos}$. We would like to minimise $\nu$ as well as $\delta$. It is difficult to simultaneously minimise both. From (20) (provided later), the value of $\delta$ is one of the factors relevant to protection against certain kinds of state guessing attacks. As explained in Remark 2 (also provided later), it is not essential to minimise $\delta$. The parameter $\nu$, on the other hand, determines the size of the maximum overlap between two inputs to the majority function corresponding to two different keystream bits. Minimising $\nu$ improves resistance to possible differential attacks (see Section 4.4). Since it is not essential to minimise $\delta$, we chose to focus on minimising $\nu$. The task of choosing ${\sf posX}$ such that $\nu$ is minimised is a combinatorial optimisation problem. There does not seem to be any good way to obtain ${\sf posX}$ such that the corresponding $\nu$ is guaranteed to be the minimum possible value. Instead, we adopted the following procedure. Given $\kappa$, $L$ and $m$, make 10000 choices of ${\sf pos}$ satisfying the constraints on $i_{1},\ldots,i_{m}$ and $j_{1},\ldots,j_{m}$, and for each choice extract ${\sf posX}$ and then compute $\nu$ from ${\sf posX}$. Finally return the ${\sf pos}$ for which the least value of $\nu$ is obtained. Then, for this ${\sf pos}$ compute the value of $\delta$. Later we provide the values of ${\sf pos}$, $\nu$ and $\delta$ for the concrete proposals that we make.
4.2 Security Analysis
The parameters $L\geq 2\kappa$, $m$, and $\mu=\lfloor\sqrt{2\kappa}\rfloor$ are the design parameters, as are the tap positions encoded by ${\sf pos}$. The parameter $\nu$ is determined by ${\sf posX}$, while the parameter $\delta$ is determined by ${\sf pos}$. The primitive connection polynomial $\tau(x)$ of the LFSR is also a parameter, as is the number of non-zero terms in $\tau(x)$. For a concrete choice targeted at a specific security level, the parameters need to be appropriately chosen so as to provide resistance against known classes of attacks. Next we provide an overview of such attacks and determine the conditions on the parameters which provide resistance to the attacks.
@@ 4.2 Security Analysis
We consider two basic attack parameters, namely the number $N$ of keystream bits that is required for a successful attack, and the time complexity $T$ of the attack. To ensure security at level $\kappa$ any attack should require $T>2^{\kappa}$. The condition $L\geq 2\kappa$ ensures that the size of the state is at least twice the size of the key. This prevents certain (theoretical) time/memory trade-off attacks [34]. We assume that at most $2^{B}$ keystream bits are generated from a single key and IV pair. So $N$ is at most $2^{B}$, as otherwise the attack cannot be mounted. Further, a basic condition is that $N\leq T$, i.e. some operation is performed on each keystream bit that is required. So for $\kappa$-bit security using $2^{B}$ keystream bits, we have $B\leq\kappa$.
Below we perform a concrete analysis of known classes of attacks so as to be able to determine conditions on the parameters which ensure resistance to such attacks. For the concrete analysis, we ignore small constants appearing in the expressions for $N$ and $T$ (i.e. we consider these constants to be 1) and focus only on the exponential components of these quantities.
@@
The degree of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1}$ is $d=2^{\lfloor\log_{2}m\rfloor}$. If $L$ is a prime, then the linear complexity of the generated keystream sequence is (${L\choose d}$ (see [49]) and so at least these many keystream bits are required to determine the linear complexity. So if
@@
These are guess-then-determine attacks. The idea is to guess some bits and then use the obtained keystream to verify the guess. The first attack of this type was described in [27] and is called the inversion attack. The inversion attack is specifically applicable to filtering functions of the form $W\oplus g(\mathbf{Z})$. Since the filtering function that we use is indeed of this form, we need to consider the resistance of the stream cipher to the inversion attack.
Suppose at some point of keystream generation, the state is $\mathbf{s}^{(0)}=(s_{L-1},\ldots,s_{0})$ and the keystream bit generated from $\mathbf{s}^{(0)}$ is $w_{0}$. From the definition of the filtering function, the bit $w_{0}$ is generated from the bits $(s_{L-1},\ldots,s_{L-2\kappa})$, and we can write $w_{0}=s_{L-2\kappa}\oplus g(s_{L-1},\ldots,s_{L-2\kappa+1})$, where $g$ is some function (defined from ${\sf MM}_{2m}$ and is non-degenerate on $2m$ variables). Recall that the tap position for $X_{1}$ is $L-1$ and the tap position for $Y_{m}$ is $L-2\kappa+1$. The gap between the tap position of $X_{1}$ and the tap position of $Y_{m}$ (including both of these tap positions) is $\chi=2\kappa-1$. Note that $s_{L-2\kappa}=w_{0}\oplus g(s_{L-1},\ldots,s_{L-2\kappa+1})$. Extending this equation in the backward direction, we obtain the following equations.
$s_{L-2\kappa}$ $=$ $w_{0}\oplus g(s_{L-1},\ldots,s_{L-2\kappa+1})$
$s_{L-2\kappa-1}$ $=$ $w_{-1}\oplus g(s_{L-2},\ldots,s_{L-2\kappa})$
$s_{L-2\kappa-2}$ $=$ $w_{-2}\oplus g(s_{L-3},\ldots,s_{L-2\kappa-1})$
$\cdots$ $\cdot$ $\cdots$
@@
The sequence $w_{0},w_{-1},w_{-2},\ldots$ is known. If we guess the values of $s_{L-1},\ldots,s_{L-2\kappa+1}$ (a total of $\chi$ bits), then we obtain $s_{L-2\kappa}$ from the first equation, using the obtained value of $s_{L-2\kappa}$ in the second equation, we obtain the value of $s_{L-2\kappa-1}$, using the obtained value of $s_{L-2\kappa-1}$ in the third equation, we obtain the value of $s_{L-2\kappa-2}$, and so on. In other words, by guessing the values of $\chi$ of the state bits, we can obtain values of $L-\chi$ previous state bits, giving us the values of all $L$ bits of the state at a point $L-\chi$ steps in the backward direction. Once the complete state is known, the original guess can be verified by generating the keystream in the forward direction and matching with the obtained keystream. The method requires $2^{\chi}$ guesses, and for each guess about $Lm$ operations are required. By our choice of the tap positions for $\mathbf{X}$ and $\mathbf{Y}$, the value of $\chi$ is $12\kappa-$ 1. So the number of guesses required to successfully mount the attack is far greater than $2^{\kappa}$. This proves the resistance of the stream cipher to the inversion attack. Note that the inversion attack is prevented due to $\chi$ being sufficiently large. This was one of the countermeasures to inversion attack that was already proposed in [27].
The inversion attack was later extended to the generalised inversion attack, the filter state guessing attack, and the generalised filter state guessing attack [27, 28, 33, 47, 58]. We briefly explain the idea behind this line of attacks. Let $n=2m+1$ be the number of variables of the filtering function. At any point of time, each of the input bits to the filtering function can be written as a linear function of the $L$ state bits of the LFSR obtained after the initialisation phase. So knowing the values of the $n$ input bits to the filtering function at any point of time provides $n$ linear equations in $L$ variables. If the values of the input bits to the filtering function at $c$ points of time are known, where $nc\geq L$, then one obtains a system of $L$ linear equations in $L$ variables and hence can solve this system to obtain the initial state of the LFSR. Of course, one does not know the values of the input bits to the filtering function at any point of time. So a guessing strategy is used.
Let as above the state of the LFSR at some point be $\mathbf{s}^{(0)}$ and the subsequent states of the LFSR be denoted as $\mathbf{s}^{(1)},\mathbf{s}^{(2)},\ldots,\mathbf{s}^{(c-1)}$. For $i=0,\ldots,c-1$, let $\mathbf{t}^{(i)}={\sf proj}(\mathbf{s}^{(i)})$ and $w_{i}=f_{2m+1}(\mathbf{t}^{(i)})$. Recall from the definition of $\delta$ that for $10\leq i<j\leq c-$ 1, $\mathbf{t}^{(i)}$ and $\mathbf{t}^{(j)}$ have an overlap of at most $\delta$ bits. So in the $c$ $n$-bit strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ there are at least $nc-\delta c(c-1)/2$ unknown bits and hence there are at least $2^{nc-\delta c(c-1)/2}$ possibilities for these $c$ strings. Since $f_{2m+1}$ is balanced, knowledge of $w_{0}$ reduces the number of possibilities for $\mathbf{t}^{(0)}$ by a factor of half, and more generally, the knowledge of $w_{0},\ldots,w_{c-1}$ reduces the number of possibilities for $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ by a factor of $2^{c}$. So at least $2^{(n-1)c-\delta c(c-1)/2}$ possibilities remain for the $c$ strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$. For each of these possibilities, we obtain a system of $L$ equations in $L$ variables, the solution of which provides the value of initial state of the LFSR. The correctness of the obtained value of the state can be determined by generating the keystream from the obtained value and matching with the actual keystream. Solving the system of $L$ linear equations requires $L^{3}$ arithmetic operations, and generating the keystream from the obtained state require about $Ln$ operations. So the complexity of the attack is $2^{(n-1)c-\delta c(c-1)/2}(L^{3}+Ln)$ operations. Suppose the following condition holds.
@@
$2^{(n-1)c-\delta c(c-1)/2}(L^{3}+Ln)$ $>$ $2^{\kappa}\mbox{ where $c$ is the least positive integer such that $nc\geq L$.}$ (20)
@@
Then no form of the state guessing attack succeeds at the $\kappa$-bit security level.
Remark 2.
In our concrete proposals given below in Table 1, the value of $c$ in (20) comes out to be 2, and so (20) becomes $2^{2(n-1)-\delta}(L^{3}+Ln)>2^{\kappa}$. Since the values of $n=2m+1$ in Table 1 are quite close to $\kappa$, ensuring $2^{2(n-1)-\delta}(L^{3}+Ln)>2^{\kappa}$ does not require $\delta$ to be too small. In fact, the values of $\delta$ (and corresponding values of $L$, $m$ and $\kappa$) given in Table 4 ensure that this inequality holds for all our proposals.
Applying the attack requires knowing the pre-image sets of 00 and $1$ of the function $f_{2m+1}$. These need to be stored separately, and depending upon the value of $w_{i}$, the appropriate pre-image set is to be used. So the storage required is $2^{2m+1}$. For the concrete proposals that we put forward later, the value of $n=2m+1$ is quite close to that of the security level $\kappa$ (see Table 1). As a result, the memory requirement will be prohitively high to apply the attack.
@@
The complexity of the above attack depends on the value of $\delta$, which is the maximum overlap of tap positions between shifts of the state vector. The recommendations for tap positions in the literature [27, 28, 33, 47, 58] aim to reduce this overlap. A commonly used recommendation is full positive difference set, i.e. the absolute values of the differences between the tap positions should be distinct. For a filtering function with $n$ variables, this recommendation results in the gap between the first and the last tap positions to be more than $21+2+\cdots+(n-1)=n(n-1)/$ 2 (since the successive differences must be distinct). So the size of the LFSR is at least quadratic in $n$. From efficiency considerations, this forces the value of $n$ to be small. By not following this recommendation, we have done away with the condition that $L$ must be at least quadratic in $n$. Instead, by considering the fundamental requirement behind state guessing attacks, we identified (20) as the condition to resist such attacks. This allows us to choose $n$ to be quite large and close to $\kappa$. Of course, this is possible due to the fact that we have an efficient method for implementing the filtering functions on such large values of $n$.
@@
Subsequent works (such as [60, 59, 39, 40]) on fast correlation attacks use vectorial decoding technique along with multiple linear approximations, use of the BKW algorithm [7], and multivariate correlation attack. The attacks are quite complex and there are no simple closed form expressions for the values of $N$ and $T$. The stream ciphers to which these attacks are applied are Grain-128a and Sosemanuk [5]. From the above discussion, we already know that the linear bias of $\mathcal{S}(257,59)$ is substantially lower than that of Grain-128a. For Sosemanuk, the best known [39] linear approximation has correlation $2^{-20.84}$, which is far greater than the linear bias $2^{-60}$ for $\mathcal{S}(257,59)$. The very low linear bias of $\mathcal{S}(257,59)$ and in general of the other stream cipher proposals in Table 1 make the attacks in [60, 59, 39, 40] inapplicable at the stated security levels.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $N^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [14]). We assume $\omega=2.8$. From Proposition 4, we assume that $a={\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})=\lceil m/2\rceil$, i.e. the algebraic immunity is equal to its lower bound. Let